# Multi-layer diffusion model of photovoltaic installations

### v2

@@ 3.3 Mean-Field Approximation
Monte Carlo computer simulations are the first choice when it comes to agent-based models. Unfortunately, as mentioned in the Introduction, they can be very time-consuming. For this reason, we derive a set of equations describing dynamics of the system (Eqs. (3)-(5)), under the assumption that each layer of the network is a complete graph of size $N\to\infty$. Thanks to this approach, we can examine a wide range of parameter values, which we could not achieve with Monte Carlo simulations in any reasonable amount of time.
@@
Figure 5: 10 simulated time trajectories of $c_{A}$ and $c_{S}$ versus numerically obtained time trajectory (from Eqs. (3)-(5)), for different values of $p$ and $a_{1}$. the AND variant (top 4) and the OR variant (bottom 4). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5a_{1}$ in all cases.
@@
Before discussing the analytical results themselves, we compare them with the simulations. For this, see Fig. 5, where we combine the latter and numerically obtained time trajectories from Eqs. (3)-(5). As shown there, mean-field gives fairly good approximation, except for the bottom row (OR variant, $p=0.2$). This is due to the fact, that in the MFA critical values of $p$ required for the system to adopt are slightly higher, as presented in the next two Figures (Figs. 6 and 7).
@@
Figure 6: Stationary states (left) and time to reach a stationary state (right) for the AND variant (top) and the OR (bottom). Left side compares numerical results (from Eqs. (3)-(5); markers) vs analytical (from Eqs. (16)-(18); continuous lines). Numerical results cover only a portion of analytical ones, as they present stationary states from a single pair of initial conditions $(c_{A}(0)=0,c_{S}(0)=0)$ only, while the latter show all the possible stationary states. Right side shows time to reach a stationary state obtained with numerical methods. Adoption probabilities $a_{1}=0.5$ (left side only) and $a_{2}=0.5a_{1}$.
@@
Mean-field time trajectories, and hence times to reach a stationary state, are obtained numerically from Eqs. (3)-(5). Analytically, we are only able to compute stationary states (Eqs. (16)-(18)). We compare the two in Fig. 6, for $a_{2}=0.5a_{1}$. Numerically computed stationary states perfectly match analytical ones. Segments not covered by the numerical results are due to the fact that the analytical solution shows all the possible stationary states, while the numerical one only those achievable from given initial conditions ($c_{A}(0)=0$, $c_{S}(0)=0$). Clearly visible here are the three possible groups of stationary states: unadopted ($c_{A}(T)\approx 0$, $c_{S}(T)\approx 0$), adopted ($c_{A}(T)\approx 1$, $c_{S}(T)\approx 1$) and disordered ($c_{A}(T)\approx 0.5$, $c_{S}(T)\approx 0.5$), although a transition between the latter two in the OR variant is very smooth. As already mentioned, the value of $a_{1}$ ($a_{1}\in(0,1]$) itself has no impact on a stationary state. Therefore, we set it arbitrarily to $0.5$ (only left side of Fig. 6). On the right, times to reach a stationary state with respect to $a_{1}$ and $p$ are presented. These drop dramatically with an increase of $a_{1}$, but only for very low values of $a_{1}$. After that, the change is unnoticeable. There are significant ,,ridges” for the values of $p$ corresponding to transitions between groups of stationary states (left side). There are two such ridges in the AND variant, but only one in the OR variant, as transition between adopted and disordered is very smooth there. These increases in times are logical, as the system needs more time to ,,decide” which path to take, and consistent with our knowledge on phase transitions.
@@
Figure 7: Stationary states (left) and time to reach a stationary state (right) for the AND variant (top) and the OR (bottom). Left side compares numerical results (from Eqs. (3)-(5); markers) vs analytical (from Eqs. (16)-(18); continuous lines). Numerical results cover only a portion of analytical ones, as they present stationary states from a single pair of initial conditions $(c_{A}(0)=0,c_{S}(0)=0)$ only, while the latter show all the possible stationary states. Right side shows time to reach a stationary state obtained with numerical methods. Adoption probabilities $a_{1}=0.5$ (left side only) and $a_{2}=0.25a_{1}$.
@@
When we decrease $a_{2}$ to $0.25a_{1}$ (see Fig. 7), transitions between adopted and disordered states become less apparent in both variants of the model. Moreover, values of $p$ needed for adoption decrease, while those needed for disorder increase, which is consistent with the simulations results.
@@ 4 Conclusions
In this research, we proposed a new agent-based model that describes both opinion formation and diffusion of photovoltaic panels. Although, it is based on the well-known model of binary opinion dynamics, the $q$-voter model, we extended it by adding a second agent attribute (adoption state) and rewrote it to a two-layer network structure. To generalize the model from a single-layer, we considered two variants. First, when the $q$-voter’s social influence is required on both layers of the network to impact one’s opinion (the AND variant). Second, where influence from just one layer is sufficient (the OR variant). We investigated the model, the effect of parameters on stationary states and times to reach them, using both Monte Carlo computer simulations and Mean-Field Approximation. For the approximation’s results, we used analytical methods wherever it was possible, and numerical ones otherwise.
@@
More specifically, within the model there exist three possible final states: unadopted system, adopted and finally, disordered one. The system transitions from the first, through the second, to the third, as the probability of independence $p$ increases. Broadly speaking, independence is initially essential for the adoption process to take off. However, as the diffusion accelerates, independence hinders innovation and ultimately prevents full adoption. Except for $a_{1}=0$, the probability of adoption has no impact on the stationary states, only on times to reach them. What impacts stationary states, however, is the relation between $a_{1}$ and probability of loosing adoption $a_{2}$. As $a_{2}$ becomes less and less of $a_{1}$, critical values of $p$ required to reach the adopted state decrease, while those required for the disordered one increase. Basically, the greater the difference between people’s willingness to install solar panels versus their willingness to get rid of them, the better for the innovation. This difference not only amplifies independence in the initial phase of adoption, but dilutes it in the final phase as well.
Regarding the studied variants, there is a significant difference between the two, AND and OR. In the AND variant, other states are achieved more easily than in the OR variant, i.e. lower values of $p$ are sufficient, but the transitions between these states are more pronounced. This is because, in the AND variant the initial negative consensus is more easily disrupted than in the OR variant, but so is the later positive consensus.
Finally, it should be noted that MFA approximates the simulation results fairly well. Although, in the simulation version the network consists of Square Lattice and two-dimensional Watts-Strogatz graph, structures pretty far from complete graph. Still, there are visible differences. Critical values of $p$ to reach various states differ slightly, but the general picture remains.
@@
In the basic research, only $c_{A}(0)=0$, $c_{S}(0)=0$ initial state was considered. However, one could investigate others as well. Here, we examine, what happens if initially a small fraction of agents possess positive opinions and adoption states, $A(0)=+1$, $S(0)=+1$. These agents could be chosen randomly or determined by some centrality measure. We compare time trajectories for the basic case (no such agents), with a number of such agents chosen randomly and the same number determined by highest degree on the second layer (degree centrality).
Figure 8: 10 simulated time trajectories of $c_{A}$ (red) and $c_{S}$ (blue) for $c_{A}(0)=0$, $c_{S}(0)=0$ (left), 100 agents with $A(0)=+1$, $S(0)=+1$ chosen randomly (middle) and 100 agents with $A(0)=+1$, $S(0)=+1$ determined by highest degree on the second layer (number of neighbors, right). The AND variant (top) and the OR variant (bottom). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5a_{1}$ in all cases.
Results are shown in Fig. 8. In AND variant there are no visible changes, as such agents easily loose either their positive opinion or adoption state when their numbers are low. Then, with the AND rule in play, they have negligible impact on the whole system. In the OR variant however, there are significant differences. Existence of such agents visibly decreases the time needed for the system to reach the adopted state. Method of their choice bears a lesser impact. This is due to the fact, that the second layer is described by the two-dimensional Watts-Strogatz graph, WS2D($N=2500$, $m=1$, $\beta=0.2$). Distribution of degrees of such a random graph corresponds to the binomial (for small networks, $N=10^{2}$) or the Poisson distribution (for larger networks, $N\geq 10^{3}$) [51]. Therefore, the network lacks agents with degrees greatly exceeding the average, who could bear an impact.
@@ Acknowledgements
T. W. was supported by the Ministry of Science and Higher Education, Poland, within the “Diamond Grant” Program through grant no. DI2019 0008 49.

### v3

@@ 3.3 Mean-Field Approximation
Monte Carlo computer simulations are the first choice when it comes to agent-based models. Unfortunately, as mentioned in the Introduction, they can be very time-consuming. For this reason, in Theorem 1 we derived a set of equations describing dynamics of the system, see Eqs. (3)-(5), under certain assumptions. Thanks to this approach, we can examine a wide range of parameter values, which we could not achieve with Monte Carlo simulations in any reasonable amount of time.
@@
Figure 5: 10 simulated time trajectories of $c_{A}$ and $c_{S}$ versus numerically obtained time trajectory from Eqs. (3)-(5), for different values of $p$ and $a_{1}$. The AND (top 4) and the OR variant (bottom 4). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5a_{1}$ in all cases.
@@
Before discussing the analytical results themselves, we compare them with the simulations. In Fig. 5, we combine the latter and numerically obtained time trajectories from Eqs. (3)-(5). As shown there, mean-field gives a fairly good approximation, except for the bottom row (OR variant, $p=0.2$). This is due to the fact, that in the MFA critical values of $p$ required for the system to adopt are slightly higher, as presented in Figs. 6 and 7.
@@
Figure 6: Stationary states (left) and time to reach a stationary state (right) for the AND (top) and the OR variant (bottom). Left side compares numerical results (from Eqs. (3)-(5); markers) vs analytical (from Eqs. (16)-(18); continuous lines). Numerical results cover only a portion of analytical ones, as they present stationary states from a single pair of initial conditions $(c_{A}(0)=0,c_{S}(0)=0)$ only, while the latter show all the possible stationary states. Right side shows time to reach a stationary state obtained with numerical methods. Adoption probabilities $a_{1}=0.5$ (left side only) and $a_{2}=0.5a_{1}$.
@@
Mean-field time trajectories, and hence times to reach a stationary state, are obtained numerically from Eqs. (3)-(5). Analytically, we are only able to compute stationary states, see Eqs. (16)-(18). We compare the two in Fig. 6, for $a_{2}=0.5a_{1}$. Numerically computed stationary states perfectly match analytical ones. Segments not covered by the numerical results are due to the fact that the analytical solution shows all the possible stationary states, while the numerical one only those achievable from given initial conditions ($c_{A}(0)=0$, $c_{S}(0)=0$). Clearly visible here are the three possible groups of stationary states: unadopted ($c_{A}(T)\approx 0$, $c_{S}(T)\approx 0$), adopted ($c_{A}(T)\approx 1$, $c_{S}(T)\approx 1$) and disordered ($c_{A}(T)\approx 0.5$, $c_{S}(T)\approx 0.5$), although the transition between the latter two in the OR variant is very smooth. As already mentioned, the value of $a_{1}\in(0,1]$ itself has no impact on a stationary state. Therefore, we set it arbitrarily to $0.5$ (only the left side of Fig. 6). On the right, times to reach a stationary state with respect to $a_{1}$ and $p$ are presented. These drop dramatically with an increase of $a_{1}$, but only for very low values of $a_{1}$. After that, the change is unnoticeable. There are significant “ridges” for the values of $p$ corresponding to transitions between groups of stationary states (left side). There are two such ridges in the AND variant, but only one in the OR variant, as the transition between adopted and disordered states is very smooth there. These increases in times are logical, as the system needs more time to “decide” which path to take, and consistent with our knowledge on phase transitions [51].
@@
Figure 7: Stationary states (left) and time to reach a stationary state (right) for the AND (top) and the OR variant (bottom). Left side compares numerical results (from Eqs. (3)-(5); markers) vs analytical (from Eqs. (16)-(18); continuous lines). Numerical results cover only a portion of analytical ones, as they present stationary states from a single pair of initial conditions $(c_{A}(0)=0,c_{S}(0)=0)$ only, while the latter show all the possible stationary states. Right side shows time to reach a stationary state obtained with numerical methods. Adoption probabilities $a_{1}=0.5$ (left side only) and $a_{2}=0.25a_{1}$.
@@
When we decrease $a_{2}$ to $0.25a_{1}$ (see Fig. 7), transitions between adopted and disordered states become less apparent in both variants of the model. Moreover, values of $p$ needed for adoption decrease, while those needed for disorder increase, which is consistent with the simulation results.
@@ 4 Conclusions
In this research, we proposed a new agent-based model that describes both opinion formation and diffusion of photovoltaic panels. Although, it is based on the well-known model of binary opinion dynamics, the $q$-voter model, we extended it by adding a second agent attribute (adoption state) and rewrote it to a two-layer network structure. To generalize the model from a single-layer, we considered two variants. First, when the $q$-voter’s social influence is required on both layers of the network to impact one’s opinion (the AND variant). Second, where influence from just one layer is sufficient (the OR variant). We investigated the model, the effect of parameters on stationary states and times to reach them, using both Monte Carlo computer simulations and the mean-field approximation. For the approximation’s results, we used analytical methods wherever it was possible, and numerical ones otherwise.
@@
More specifically, within the model there exist three possible final states: unadopted system, adopted and finally, disordered one. The system transitions from the first, through the second, to the third, as the probability of independence $p$ increases. Broadly speaking, independence is initially essential for the adoption process to take off. However, as the diffusion accelerates, independence hinders innovation and ultimately prevents full adoption. Except for $a_{1}=0$, the probability of adoption has no impact on the stationary states, only on times to reach them. What impacts stationary states, however, is the relation between $a_{1}$ and probability of unadopting $a_{2}$. As $a_{2}$ becomes much smaller than $a_{1}$, critical values of $p$ required to reach the adopted state decrease, while those required for the disordered one increase. Basically, the greater the difference between people’s willingness to install solar panels versus their willingness to get rid of them, the better for the innovation. This difference not only amplifies independence in the initial phase of adoption, but dilutes it in the final phase as well.
Regarding the studied variants, there is a significant difference between the two, AND and OR. In the AND variant, adopted and disordered states are achieved more easily than in the OR variant, i.e. lower values of $p$ are sufficient, but the transitions between these states are more pronounced. This is because, in the AND variant the initial negative consensus is more easily disrupted than in the OR variant, but so is the later positive consensus.
Finally, it should be noted that MFA approximates the simulation results fairly well. Despite the fact that in the simulation version the network consists of a Square Lattice and a two-dimensional Watts-Strogatz graph, structures pretty different from complete graph. Still, there are visible differences. Critical values of $p$ to reach various states differ slightly, but the general picture remains.
@@
In the results presented above, only the $c_{A}(0)=0$, $c_{S}(0)=0$ initial state was considered. However, one could investigate others as well. Here, we examine, what happens if initially a small fraction of agents possess positive opinions and adoption states, $A(0)=+1$, $S(0)=+1$. These agents could be chosen randomly or determined by some centrality measure. We compare time trajectories for the basic case (no such agents), with the number of such agents chosen randomly and the same number determined by the highest degree on the second layer (degree centrality).
Figure 8: 10 simulated time trajectories of $c_{A}$ (red) and $c_{S}$ (blue) for $c_{A}(0)=0$, $c_{S}(0)=0$ (left), 100 agents with $A(0)=+1$, $S(0)=+1$ chosen randomly (middle) and 100 agents with $A(0)=+1$, $S(0)=+1$ determined by the highest degree on the second layer (number of neighbors, right). The AND (top) and the OR variant (bottom). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5a_{1}$ in all cases.
Results are shown in Fig. 8. In the AND variant there are no visible changes, as such agents easily loose either their positive opinion or adoption state when their numbers are low. Then, with the AND rule in play, they have negligible impact on the whole system. In the OR variant however, there are significant differences. Existence of such agents visibly decreases the time needed for the system to reach the adopted state. The method of their choice bears a lesser impact. This is due to the fact, that the second layer is the two-dimensional Watts-Strogatz graph, WS2D($N=2500$, $m=1$, $\beta=0.2$). Degree distribution of such a random graph corresponds to the binomial (for small networks, $N=10^{2}$) or the Poisson distribution (for larger networks, $N\geq 10^{3}$) [52]. Therefore, the network lacks agents with degrees greatly exceeding the average, who could bear an impact.
@@ Acknowledgements
This study was supported by the Ministry of Science and Higher Education, Poland, within the “Diamond Grant” Program through grant no. DI2019 0008 49.